Convert 9.29×10−3 m to the equivalent length in nanometers.
1
Expert's answer
2019-09-09T04:57:16-0400



Task:

Convert 9.29×10−3 m to the equivalent length in nanometers


Solution:

1 Meter = 1×109 Nanometers.

Then,

[9.29×10−3 meter * 1×109 Nanometers] / 1 meter = 9290000 Nanometers = 9290000 nm


Answer:  9.29×10−3 m = 9290000 nm = 9.29×106